# Depth-Weighted Detection of Behaviours of Risk in People with Dementia using Cameras

### v2

@@ 3.5 Behaviours of Risk Detection
Both F1-score and MCC are threshold-based metrics; hence they evaluate performance over a single threshold. The area under the curve of receiver operating characteristic (AUROC) evaluates the performance over TPR and FPR over a range of thresholds. Hence, the primary metrics employed for evaluating the performance of methods in this work are MCC, F1-score and AUROC. We also report additional metrics for reference, namely, TPR, true negative rate (TNR), FPR, FNR, gmean ($\sqrt{TPR*TNR}$), precision, recall, specificity, and balanced accuracy [30]. The evaluation metrics use reconstruction error as the score to identify behaviours of risk as anomalies during testing.
@@
TPR 0.945 0.950 0.641 0.939 0.773 0.552
TNR 0.628 0.581 0.713 0.684 0.742 0.780
FPR 0.372 0.419 0.287 0.316 0.258 0.220
FNR 0.055 0.050 0.359 0.061 0.227 0.448
Gmean 0.770 0.743 0.676 0.801 0.758 0.657
Precision 0.165 0.150 0.148 0.187 0.189 0.163
Recall 0.945 0.950 0.641 0.939 0.773 0.552
Specificity 0.628 0.581 0.713 0.684 0.742 0.780
Balanced Accuracy 0.786 0.765 0.677 0.811 0.758 0.666
MCC 0.301 0.276 0.198 0.335 0.293 0.200
F1-score 0.281 0.259 0.240 0.313 0.304 0.252
AUROC 0.844 0.779 0.754 0.852 0.823 0.784
@@
TPR 0.949 0.728 0.502 0.844 0.412 0.751
TNR 0.611 0.668 0.812 0.686 0.832 0.719
FPR 0.389 0.332 0.188 0.314 0.168 0.281
FNR 0.051 0.272 0.498 0.156 0.588 0.249
Gmean 0.761 0.697 0.638 0.761 0.586 0.735
Precision 0.162 0.148 0.174 0.176 0.163 0.175
Recall 0.949 0.728 0.502 0.844 0.412 0.751
Specificity 0.611 0.668 0.812 0.686 0.832 0.719
Balanced Accuracy 0.780 0.698 0.657 0.765 0.622 0.735
MCC 0.290 0.215 0.200 0.290 0.164 0.264
F1-score 0.276 0.246 0.259 0.291 0.234 0.284
AUROC 0.810 0.766 0.790 0.810 0.757 0.792
@@
Gmean 0.698 0.647 0.455 0.722 0.680 0.355
Precision 0.118 0.111 0.139 0.149 0.119 0.119
Recall 0.919 0.649 0.230 0.730 0.743 0.135
Specificity 0.530 0.645 0.903 0.714 0.622 0.931
Balanced Accuracy 0.724 0.647 0.566 0.722 0.682 0.533
@@ 4 Results and Discussion
In this section, we evaluate the performance of the proposed approach for FP reduction and detecting behaviours of risk in PwD. We conduct an ablation analysis of different components of the proposed approach and the effect of frame size and frame rate on the performance of detecting behaviours of risk. We further perform cross-camera, participant and sex-specific analysis.
Performance Evaluation: We investigate the effectiveness of the proposed approach for reducing FP and detecting behaviours of risk by evaluating the improvement in performance after integrating the proposed approach into existing methods. As discussed in Section 3.5, depth variants were created after integrating the proposed approach into the existing methods. Tables 6, 6 and 6 compare the performance of the depth variants with their existing method counterparts for behaviours of risk detection for Cam1, Cam2 and Cam3, respectively. The best values for FPR and primary evaluation metrics, MCC, F1-score and AUROC, are marked in bold. It can be observed that depCAE performed mostly better than 3DCAE in terms of MCC, F1-score and AUROC for all three cameras. depCTBAE performed better than CTBAE for Cam1 and Cam3, while depVAE performed better than VAE for Cam1 and Cam2. depCAE recorded lower FPR than 3DCAE for all three cameras. depCTBAE and depVAE recorded lower FPR than CTBAE and VAE, respectively, for two out of three cameras. The methods performed better for Cam1 and Cam2 in comparison to Cam3, which can be attributed to the major portion of Cam3 view being occluded by the wall. This made it difficult to gauge the behaviours of risk events clearly in Cam3. For the case of secondary evaluation metrics, gmean, precision and specificity, the depth variants mostly performed better than the existing methods across the three cameras. The depCAE method achieved the best MCC, F1-score and AUROC overall for all three cameras; hence further analysis in this section is performed using depCAE method. The depCAE method achieved a MCC, F1-score and AUROC of $0.335$, $0.313$, $0.852$ for Cam1, $0.29$, $0.291$, $0.81$ for Cam2, and $0.235$, $0.248$, $0.768$ for Cam3, respectively. While depVAE gave lower FPR as compared to other methods, it gave significantly higher FNR. In such cases, it is important to consider a balanced evaluation over both FPR and FNR, which is provided by F1-score, where depCAE performed the best. It can be inferred from the observations that the proposed approach of weighing the loss with pixel depth is indeed effective in reducing false positives and improving detection performance. Further, the proposed approach can be easily integrated into any existing autoencoder-based anomaly detection method.
@@
Gmean 0.760 0.805 0.801 0.743 0.791 0.761 0.624 0.647 0.722
Precision 0.158 0.202 0.187 0.158 0.188 0.176 0.098 0.135 0.149
Recall 0.967 0.895 0.939 0.879 0.907 0.844 0.905 0.554 0.730
Specificity 0.598 0.725 0.684 0.628 0.690 0.686 0.430 0.756 0.714
Balanced Accuracy 0.782 0.81 0.811 0.754 0.798 0.765 0.668 0.655 0.722
@@
Gmean 0.801 0.808 0.761 0.645 0.722 0.653
Precision 0.187 0.199 0.176 0.198 0.149 0.105
Recall 0.939 0.917 0.844 0.494 0.730 0.851
Specificity 0.684 0.712 0.686 0.841 0.714 0.501
Balanced Accuracy 0.811 0.815 0.765 0.668 0.722 0.676
@@
Gmean 0.801 0.750 0.761 0.704
Precision 0.187 0.162 0.176 0.139
Recall 0.939 0.862 0.844 0.848
Specificity 0.684 0.653 0.686 0.585
Balanced Accuracy 0.811 0.757 0.765 0.717
@@
Table 11: depCAE performance stratified over sex.
Cam1 Cam2 Cam3
F1-score AUROC F1-score AUROC F1-score AUROC
Male (N=2) 0.310 0.837 0.266 0.738 0.160 0.637
Female (N=4) 0.313 0.856 0.294 0.817 0.276 0.806
Sex-Specific Analysis: It is important to investigate the performance of depCAE across population groups. Table 11 presents the performance of depCAE stratified based on sex across three cameras. As can be seen in Table 3, some participants did not show any behaviours of risk event in one or two cameras, leading to the situation with only 6 participants with differing participant ids for any individual camera. However, the overall sex-specific analysis was done for all 9 participants. A similar behaviour of risk detection performance was observed for both males and females for Cam1. However, for cameras $2$ and $3$ it was observed that the depCAE method performed better for females than males. The reason for this can be attributed to the presence of more females than males in the study, leading to the skewed class distribution due to the presence of more behaviours of risk windows for females than males, enhancing the method’s learning and generalization capabilities for females.
@@ 5 Conclusion and Future Work
Caring for individuals with dementia in care settings is increasingly challenging due to growing patient numbers and understaffing issues. The behavioural and psychological symptoms of dementia can lead to incidents that jeopardize the health and safety of patients, staff, and caregivers. Leveraging existing video infrastructure can enable the development of innovative deep-learning approaches to detect such behaviours, thereby preventing injuries and enhancing patient care. As a step in this direction, we proposed an approach that uses a novel depth-weighted loss to enforce equivalent importance to both near and far events and uses training outliers to determine the anomaly threshold. The depCAE achieved AUROC of $0.852$, $0.81$ and $0.768$ for the three cameras, outperforming other methods. The proposed approach recorded lower FPR than the existing methods across all three cameras, thus reducing the false alarms in detecting behaviours of risk in PwD. Further, the proposed approach consistently demonstrated its effectiveness and can easily be integrated into any existing autoencoder-based anomaly detection method. This motivates further multidisciplinary research for the deployment of video surveillance-based behaviours of risk detection systems in nursing and care settings for PwD. Similar multidisciplinary approaches have shown to be helpful in medical settings [31, 32]. A video-based behaviours of risk detection system will help to improve the quality of life for the residents in care homes.
Future work can involve validating our approach on data collected from an LTC that typically encounters fewer behaviours of risk incidents. The influence of camera placement on the global system performance needs to be further investigated by evaluating how the position, angle, and orientation of the cameras impact the system’s ability to accurately and reliably detect behaviours of risk. Further, the effect of utilizing the depth map as part of data as an additional input channel can be investigated. 3D object reconstruction-based approaches [33, 34] can be used to circumvent the problem of occlusions and improve the performance of detecting behaviours of risk as anomalies. Pose tracking cameras that directly extract body joints information without recording actual scene [35] can be employed in private areas to detect behaviours of risk in a privacy-protecting manner [4, 36]. It can be a time-consuming and expensive task to train a method separately on each new scene or environment. A lot of time and resources can be saved if a method trained on one scene can generalize well in detecting behaviours of risk in an entirely new scene. Fine-tuning on target camera using few-shot learning and transfer learning can be employed to improve the performance of depCAE for cross-camera generalization.

### v3

@@ 3.4 Behaviours of Risk Detection
Evaluation Metrics: The proposed approach aims to reduce false positives; however, it is also important to consider false negatives when detecting behaviours of risk in PwD. Hence, we consider a metric, i.e. F1-score, that can give a balanced estimate of the performance on both false positive rate (FPR) and false negative rate (FNR). F1-score is the harmonic mean of precision and recall, and provides a reasonable performance estimate over true positive rate (TPR), FPR and FNR, simultaneously. Further, Matthew’s Correlation Coefficient (MCC) gives an overall performance estimate over true positives (TP), true negatives (TN), false positives (FP) and false negatives (FN) and is expressed as, $MCC=(TP*TN-FP*FN)/\sqrt{(TP+FP)*P*N*(TN+FN)}$. Both F1-score and MCC are threshold-based metrics; hence they evaluate performance over a single threshold. The area under the curve of receiver operating characteristic (AUROC) evaluates the performance over TPR and FPR over a range of thresholds. Hence, the primary metrics employed for evaluating the performance of methods in this work are MCC, F1-score and AUROC. We also report additional metrics for reference, namely, TPR, true negative rate (TNR), FPR, and FNR. The performance is evaluated using depth-weighted loss as the score to identify behaviours of risk during testing.
@@
TPR 0.945 0.950 0.641 0.939 0.79 0.541
TNR 0.628 0.581 0.713 0.684 0.747 0.777
FPR 0.372 0.419 0.287 0.316 0.253 0.223
FNR 0.055 0.050 0.359 0.061 0.21 0.459
MCC 0.301 0.276 0.198 0.335 0.305 0.191
F1-score 0.281 0.259 0.240 0.313 0.313 0.246
AUROC 0.844 0.779 0.754 0.852 0.821 0.784
@@
TPR 0.949 0.728 0.502 0.844 0.444 0.751
TNR 0.611 0.668 0.812 0.686 0.819 0.719
FPR 0.389 0.332 0.188 0.314 0.181 0.281
FNR 0.051 0.272 0.498 0.156 0.556 0.249
MCC 0.290 0.215 0.200 0.290 0.172 0.264
F1-score 0.276 0.246 0.259 0.291 0.238 0.284
AUROC 0.810 0.766 0.790 0.810 0.771 0.792
@@ 4 Results and Discussion
In this section, we evaluate the performance of the proposed approach for FP reduction and detecting behaviours of risk in PwD. We conduct an ablation analysis of different components of the proposed approach and the effect of frame size and frame rate on the performance of detecting behaviours of risk. We further perform cross-camera, and participant-specific analysis.
Performance Evaluation: We investigate the effectiveness of the proposed approach for reducing FP and detecting behaviours of risk by evaluating the improvement in performance after integrating the proposed approach into existing methods. As discussed in Section 3.4, depth variants were created after integrating the proposed approach into the existing methods. Tables 6, 6 and 6 compare the performance of the depth variants with their existing method counterparts for behaviours of risk detection for Cam1, Cam2 and Cam3, respectively. The best values for FPR and primary evaluation metrics, MCC, F1-score and AUROC, between each method and its corresponding depth variant are marked in bold. It can be observed that depCAE performed mostly better than 3DCAE in terms of MCC, F1-score and AUROC for all three cameras. depCTBAE performed better than CTBAE for Cam1 and Cam3, while depVAE performed better than VAE for Cam1 and Cam2. depCAE recorded lower FPR than 3DCAE for all three cameras. depCTBAE and depVAE recorded lower FPR than CTBAE and VAE, respectively, for two out of three cameras. The methods performed better for Cam1 and Cam2 in comparison to Cam3, which can be attributed to the major portion of Cam3 view being occluded by the wall. This made it difficult to gauge the behaviours of risk events clearly in Cam3. While depVAE gave lower FPR as compared to other methods, it gave significantly higher FNR. In such cases, it is important to consider a balanced evaluation over both FPR and FNR, which is provided by F1-score, where depCAE performed the best. It can be inferred from the observations that the proposed approach of weighing the loss with pixel depth is indeed effective in reducing false positives and improving detection performance. Further, the proposed approach can be easily applied into any existing autoencoder-based anomaly detection method. The depCAE method achieved a MCC, F1-score and AUROC of $0.335$, $0.313$, $0.852$ for Cam1, $0.29$, $0.291$, $0.81$ for Cam2, and $0.235$, $0.248$, $0.768$ for Cam3, respectively. The depCAE method achieved the best MCC, F1-score and AUROC overall for all three cameras; hence further analysis in this section is performed using depCAE method.
@@ 5 Conclusion and Future Work
We proposed an approach that uses a novel depth-weighted loss to enforce equivalent importance to both near and far events and uses training outliers to determine the anomaly threshold. The proposed approach, when applied to existing methods, recorded higher F1-score, MCC and AUROC while keeping lower FPR than the existing methods across all three cameras, thus performing better while reducing the false alarms in detecting behaviours of risk in PwD. The proposed approach consistently demonstrated its effectiveness and can easily be applied into any existing autoencoder-based anomaly detection method. This motivates further multidisciplinary research for the deployment of video surveillance-based behaviours of risk detection systems in nursing and care settings for PwD. Similar multidisciplinary approaches have shown to be helpful in medical settings [26, 27].
Future work can involve validating our approach on data collected from an LTC that typically encounters fewer behaviours of risk incidents. Further, the effect of utilizing the depth map as part of data as an additional input channel can be investigated. Pose tracking cameras that directly extract body joints information without recording actual scene [28] can be employed in private areas to detect behaviours of risk in a privacy-protecting manner [4, 29]. It can be a time-consuming and expensive task to train a method separately on each new camera scene or environment. Fine-tuning on target camera using few-shot learning and transfer learning can be employed to improve the performance of depCAE for cross-camera generalization.